# Describing Functions and Phase Response Curves of Excitable Systems

### v1

@@ II-A Describing functions
The describing function method was originally developed for the feedback interconnection of a linear time-invariant system with a static nonlinearity. The describing function $N(A,\omega)$ of the static nonlinearity expresses the gain and phase shift introduced by the nonlinear element when driven by a sinusoidal input of amplitude $A$ and frequency $\omega$:
@@ IV-B Predicting ring network oscillations
A straightforward application of the eDF is in predicting network oscillation existence and period: A ring network of $N$ excitable systems permits a rhythm with period $T$ if the relative phases of its components sum to $1$.
@@
We can now compare the predictions of the proposed framework to the behavior of a simulated neuronal network. As it turns out, the prediction of oscillation period for an I-I motif (the HCO), is quite accurate, as shown in Table I. The mismatch between the prediction and the realistic oscillation period stems from the modeling mismatch between the Dirac delta impulses and the spiking waveform. Since the spike width somewhat depends on its cause (excitation or rebound) in the HH-model, so does the exact trajectory of the synaptic current. For example, rebound spikes tend to have a slightly longer duration, thus our input impulses, which have a constant width, may underestimate the onset time. Such a negative bias on the eDF could be mitigated by balancing out this effect by increasing the synaptic conductance, or by increasing the effective impulse duration. Because of such underestimation, the prediction may become more inaccurate in the nonlinear regime of small periods.

### v2

@@ II-A Describing functions
The describing function method was originally developed for the feedback interconnection of a linear time-invariant system with a static nonlinearity. The describing function $N(A,\omega)$ of the static nonlinearity expresses the gain and phase shift introduced by the nonlinear element when driven by a sinusoidal input of amplitude AA and frequency $\omega$:
@@ IV-B Predicting ring network oscillations
A straightforward application of the eDF is in predicting network oscillation existence and period: A ring network of NN excitable systems permits a rhythm with period $T$ if the relative phases of its components sum to 11.
@@
We can now compare the predictions of the proposed framework to the behavior of a simulated neuronal network. As it turns out, the prediction of oscillation period for an I-I motif (the HCO), is quite accurate, as shown in Table I. The mismatch between the prediction and the realistic oscillation period stems from the modeling mismatch between the Dirac delta impulses and the spiking waveform. Since the spike width somewhat depends on its cause (excitation or rebound) in the $H$-model, so does the exact trajectory of the synaptic current. For example, rebound spikes tend to have a slightly longer duration, thus our input impulses, which have a constant width, may underestimate the onset time. Such a negative bias on the eDF could be mitigated by balancing out this effect by increasing the synaptic conductance, or by increasing the effective impulse duration. Because of such underestimation, the prediction may become more inaccurate in the nonlinear regime of small periods.